Aspire Faculty ID #3749 · Topic: NIMCET 2018 · Just now
NIMCET 2018

A and B play a game where each is asked to select a number from 1 to 25. If the two number match, both of them win a prize. The probability that they will not win a prize in a single trial is :

Solution

The total number of ways in which numbers can be choosed =25 x 25=625  The number of ways in which either players can choose same numbers = 25  
Probability that they win a prize = 25/625 = 1/25 
Thus, the probability that they will not win a prize in a single trial = 1 - 1/25 = 24/25

Previous 10 Questions — NIMCET 2018

Nearest first

Next 10 Questions — NIMCET 2018

Ascending by ID
Ask Your Question or Put Your Review.

loading...